# Memory Is Not the Bottleneck: Cost-Efficient Continual Learning via Weight Space Consolidation

### v3

@@ A.7 Societal Impacts
A.8 Assets
In our work, many existing assets were used. For the implementation of the models and the learning algorithms, we have used the Pytorch [52] (BSD-3 license) and Huggingface [65] libraries (varies on each library. For instance, the core Transformers library uses an Apache 2.0 license). All of the datasets we have used are public datasets. For instance: CIFAR-100 (MIT License), ImageNet-100 (Custom, Non-commercial Research Only, has a separate terms of use). Regarding the language datasets in TRACE: C-STANCE (CC BY-NC 4.0) ), FOMC (CC BY-NC-SA 4.0), MeetingBank (CC BY-NC-SA 4.0), Py150 (MIT License), ScienceQA (MIT License), NumGLUE-cm (MIT License), NumGLUE-ds (MIT License), and 20Minuten (CC BY-NC-SA 4.0).

### v5

@@ A.8 Societal Impacts
A.9 Assets
In our work, many existing assets were used. For the implementation of the models and the learning algorithms, we have used the Pytorch (Paszke et al., 2019) (BSD-3 license) and Huggingface (Wolf et al., 2020) libraries (varies on each library. For instance, the core Transformers library uses an Apache 2.0 license). All of the datasets we have used are public datasets. For instance: CIFAR-100 (MIT License), ImageNet-100 (Custom, Non-commercial Research Only, has a separate terms of use). Regarding the language datasets in TRACE: C-STANCE (CC BY-NC 4.0) ), FOMC (CC BY-NC-SA 4.0), MeetingBank (CC BY-NC-SA 4.0), Py150 (MIT License), ScienceQA (MIT License), NumGLUE-cm (MIT License), NumGLUE-ds (MIT License), and 20Minuten (CC BY-NC-SA 4.0).